Heavy Metal Corporation is expected to generate the following free cash flows over the next five​ years:

Year

1

2

3

4

5

FCF​ ($ million)

51.9

67.7

76.9

73.3

83.1

​Thereafter, the free cash flows are expected to grow at the industry average of

4.1 %4.1%

per year. Using the discounted free cash flow model and a weighted average cost of capital of

13.5 %13.5%​:

a.  Estimate the enterprise value of Heavy Metal.

b.  If Heavy Metal has no excess​ cash, debt of

$ 297$297

​million, and

3939

million shares​ outstanding, estimate its share price.

a.  Estimate the enterprise value of Heavy Metal.
Answer #1

a). Enterprise Value = [FCF1 / (1 + WACC)] + [FCF2 / (1 + WACC)2] + [FCF3 / (1 + WACC)3] + [FCF4 / (1 + WACC)4] + [FCF5 / (1 + WACC)5] + [{FCF5 * (1 + g)} / {(WACC - g) * (1 + WACC)5}]

= [51.9 / (1 + 0.135)] + [67.7 / (1 + 0.135)2] + [76.9 / (1 + 0.135)3] + [73.3 / (1 + 0.135)4] + [83.1 / (1 + 0.135)5] + [{83.1 * (1 + 0.041)} / {(0.135 - 0.041) * (1 + 0.135)5}]

= 45.73 + 52.55 + 52.59 + 44.17 + 44.12 + 488.59 = $727.75 million

b). Equity Value = Enterprise Value - Debt Value

= 727.75 - 297 = $430.75 million

Share Price = Equity Value / Shares Outstanding = $430.75 million / 39 million = $11.04